Grade 11Differential Calculus

Prove that the derivative of (sec x + tan x)/(sec x - tan x) is 2cos x/(1 - sin x)2

Arjya Dutta
7 Years agoGrade 11

1 Answer

Arun
7 Years ago
Dear student
 
we can write it as 
 
(sec x + tan x)2
 
hence its derivative =
 
2 (sec x + tan x) * (sec x tan x + sec2 x)
 
 = 2 (1 + sinx )( 1 + sin x) / cos3x
 
 =  2 (1 + sin x) / cos x ( 1- sinx)
 
 = 2 (1 – sin2 x) / cos x ( 1 – sin x)2
 
 = 2 cos x / (1- sin x)2
 
hope it helps